# Consistent Opponent Modeling of Static Opponents in Imperfect-Information Games

### v4

@@ Abstract
The goal of agents in multi-agent environments is to maximize total reward against the opposing agents that are encountered. Following a game-theoretic solution concept, such as Nash equilibrium, may obtain a strong performance in some settings; however, such approaches fail to capitalize on historical and observed data from repeated interactions against our opponents. Opponent modeling algorithms integrate machine learning techniques to exploit suboptimal opponents utilizing available data; however, the effectiveness of such approaches in imperfect-information games to date is quite limited. We show that existing opponent modeling approaches fail to satisfy a simple desirable property even against static opponents drawn from a known prior distribution; namely, they do not guarantee that the model approaches the opponent’s true strategy even in the limit as the number of game iterations approaches infinity. We develop a new algorithm that is able to achieve this property and runs efficiently by solving a convex minimization problem based on the sequence-form game representation using projected gradient descent. The algorithm is guaranteed to efficiently converge to the opponent’s true strategy given observations from gameplay and possibly additional historical data if it is available.
@@ 1 Introduction
In perfect-information games, opponent modeling is relatively straightforward. First consider the case of simultaneous-move games, often represented using the strategic form. Let $S_{i}$ denote the set of pure strategies for player $i$, $S$ denote the set of strategy profiles (vectors of strategies for each player), and $u_{i}:S\rightarrow\mathbf{R}$ the utility function for player $i$. For now suppose the game has two players and that we play the role of player 1. In the absence of any historical data on our opponent, we can use a prior distribution that is Dirichlet with mean on the mixed strategy that plays each pure strategy with equal probability. This can be done by initializing $m^{0}_{s_{2,j}}=\alpha$ for all $s_{2,j}\in S_{2},$ where $\alpha$ is a positive constant and $S_{2}$ is the pure strategy space of the opponent. At each iteration $t$, we assume that the opponent will play pure strategy $s_{2,j}$ with probability $\frac{m^{t}_{s_{2,j}}}{\sum_{k}m^{t}_{s_{2,k}}},$ and we can simply play any best response to this strategy. If the opponent plays $s_{2,j}$ at time $t$, then we set $m^{t}_{s_{2,j}}=m^{t-1}_{s_{2,j}}+1$, and otherwise we set $m^{t}_{s_{2,j}}=m^{t-1}_{s_{2,j}}.$ Then our new opponent model will again be the mean of the new Dirichlet posterior distribution given our prior and observations. By the law of large numbers, the opponent model will approach the opponent’s true static mixed strategy as the number of game iterations $t$ approaches infinity. If historical data is available, then we could use a Dirichlet prior with a different mean strategy by modifying the initial values $m^{0}_{s_{2,j}}$ accordingly. Thus, in two-player strategic-form games, we can straightforwardly perform efficient opponent modeling and exploitation against static mixed strategies. This approach can also be straightforwardly extended to apply to more than two players whose strategies are played independently. In perfect-information sequential games, opponent modeling is also straightforward. Such games are typically represented as trees. We can simply apply the same procedure as above independently at each decision node of the tree. This also applies whether or not chance moves are allowed in the game tree.
@@
To perform opponent modeling we assume that we are in the repeated game setting where the game is repeated for some number of iterations. This number may be finite, infinite, or unknown (and potentially infinite). Defining repeated games in the strategic-form setting is relatively straightforward. At each iteration $t$, we know our own mixed strategy $\sigma^{t}_{i},$ as well as the vector of pure strategies selected by all players $\mathbf{s^{t}}$ (though of course we do not know opponents’ mixed strategies). We remember all of these strategies for all subsequent iterations. Therefore, at iteration $t$, our strategy can be dependent on $\sigma^{1}_{i},\ldots,\sigma^{t-1}_{i}$ as well as $\mathbf{s^{1}},\ldots,\mathbf{s^{t-1}}$. In perfect-information sequential games, we know our own mixed strategy $\sigma^{t}_{i},$ as well as the actions taken by all players at nodes along the path of play (we do not know what an opposing player would have done at a node that was not reached). Again, at time $t$ our strategy can be dependent on these values for times from 1 to t1.t-1.
@@
Now that we have defined the repeated game setting for imperfect-information games, we can consider opponent modeling algorithms for this setting. It turns out that the general goal is still the same; given a prior distribution (possibly an uninformative Dirichlet distribution whose mean selects each available action with equal probability) and our observations made during game play, we would like our model for the opponent’s strategy to be the mean of the posterior distribution. This is justified by the following theorem and corollary [3]. Theorem 1 shows that the expected payoff of a mixed strategy $\sigma_{i}$ against a probability distribution with pdf $f_{-i}$ over mixed strategies equals the expected payoff of playing $\sigma_{i}$ against the mean $\overline{f_{-i}}$ of $f_{-i}$. This applies to both strategic-form and extensive-form games (with perfect and imperfect information) as well as to any number of players ($\sigma_{-i}$ can be a joint strategy profile for all opposing agents). From this result Corollary 1 immediately follows if we let $f_{-i}$ be the posterior probability distribution for the opponent’s strategy $\sigma_{-i}$ given observations $x$, denoted as $p(\sigma_{-i}|x)$.
@@ 2 Consistent opponent modeling
If we can show that BBR is not consistent in a strategic-form game, this directly implies that it is also not consistent in extensive-form games of perfect and imperfect information. Consider the game of Rock-Paper-Scissors and suppose that $\sigma^{*}_{-i}=(0.8,0.1,0.1)$ for (R,P,S). Suppose that we are using BBR with $k=3$ sampled strategies: (0.5,0.3,0.2), (0.3,0.5,0.2), (0.2,0.3,0.5). BBR chooses the opponent model to play each of the sampled strategies in proportion to their posterior value; therefore, the opponent model is always a convex combination of the sampled strategies. Since $\sigma^{*}_{-i}$ is clearly not close to any convex combination of the sampled strategies, clearly $M_{t}$ cannot approach $\sigma^{*}_{-i}$. ∎
As the proof of Proposition 1 shows, clearly BBR is constrained to only selecting opponent models that are convex combinations of the sampled strategies. So if the true strategy $\sigma^{*}_{-i}$ is outside the convex hull of the sampled strategies, BBR has no hope of the opponent model approaching the true strategy even if the game is repeated infinitely many times. Of course, in a strategic-form game such as Rock-Paper-Scissors we could include all pure strategies of the game in our set of “samples,” which would ensure that every mixed strategy is within the convex hull of the samples. However, in extensive-form games there are exponentially many pure strategies in the size of the game tree and it is not feasible to include all of them. In fact, Proposition 2 shows that BBR may not be consistent even if $\sigma^{*}_{-i}$ is in the convex hull of the sampled strategies.
@@ 3 Algorithm
Proposition 4.
The algorithm is consistent.
@@
At each iteration $t$ the algorithm computes the modeled strategy for the opponent $M_{t}$ that maximizes the posterior distribution. In the limit as the number of iterations goes to infinity the posterior distribution approaches a point mass around the true value $\sigma^{*}_{-i}.$ So $\lim_{t\rightarrow\infty}M_{t}=\sigma^{*}_{-i}.$
@@ 6 Conclusion
We introduced a new algorithm that computes the strategy for the opponent that maximizes the posterior distribution assuming independent Dirichlet prior distributions at each information set. The Dirichlet prior is very common for modeling vectors in a multidimensional simplex. With no further information available, we could choose a Dirichlet distribution with equal parameters for each action (whose mean selects each action with equal probability). If historical information is available, we could alternatively use a Dirichlet prior distribution whose mean strategy is calculated from the data. An additional option that has proven successful in certain applications is to use a Dirichlet prior whose mean strategy is a Nash equilibrium (or approximation of one) [2, 4]. Our algorithm runs efficiently by solving a convex minimization problem using projected gradient descent. We show that our algorithm achieves consistency, unlike several prior algorithms that utilize sampling.

### v5

# Consistent Opponent Modeling in Imperfect-Information Games

@@ Abstract
The goal of agents in multi-agent environments is to maximize total reward against the opposing agents that are encountered. Following a game-theoretic solution concept, such as Nash equilibrium, may obtain a strong performance in some settings; however, such approaches fail to capitalize on historical and observed data from repeated interactions against our opponents. Opponent modeling algorithms integrate machine learning techniques to exploit suboptimal opponents utilizing available data; however, the effectiveness of such approaches in imperfect-information games to date is quite limited. We show that existing opponent modeling approaches fail to satisfy a simple desirable property even against static opponents drawn from a known prior distribution; namely, they do not guarantee that the model approaches the opponent’s true strategy even in the limit as the number of game iterations approaches infinity. We develop a new algorithm that is able to achieve this property and runs efficiently by solving a convex minimization problem based on the sequence-form game representation using projected gradient descent. The algorithm is guaranteed to efficiently converge to the opponent’s true strategy under standard Bayesian identifiability and visitation assumptions, given observations from gameplay and possibly additional historical data if it is available.
@@ 1 Introduction
In perfect-information games, opponent modeling is relatively straightforward. First consider the case of simultaneous-move games, often represented using the normal form. Let $S_{i}$ denote the set of pure strategies for player $i$, $S$ denote the set of strategy profiles (vectors of strategies for each player), and $u_{i}:S\rightarrow\mathbf{R}$ the utility function for player $i$. For now suppose the game has two players and that we play the role of player 1. In the absence of any historical data on our opponent, we can use a prior distribution that is Dirichlet with mean on the mixed strategy that plays each pure strategy with equal probability. This can be done by initializing $m^{0}_{s_{2,j}}=\alpha$ for all $s_{2,j}\in S_{2},$ where $\alpha$ is a positive constant and $S_{2}$ is the pure strategy space of the opponent. At each iteration $t$, we assume that the opponent will play pure strategy $s_{2,j}$ with probability $\frac{m^{t}_{s_{2,j}}}{\sum_{k}m^{t}_{s_{2,k}}},$ and we can simply play any best response to this strategy. If the opponent plays $s_{2,j}$ at time $t$, then we set $m^{t}_{s_{2,j}}=m^{t-1}_{s_{2,j}}+1$, and otherwise we set $m^{t}_{s_{2,j}}=m^{t-1}_{s_{2,j}}.$ Then our new opponent model will again be the mean of the new Dirichlet posterior distribution given our prior and observations. By the law of large numbers, the opponent model will approach the opponent’s true static mixed strategy as the number of game iterations $t$ approaches infinity. If historical data is available, then we could use a Dirichlet prior with a different mean strategy by modifying the initial values $m^{0}_{s_{2,j}}$ accordingly. Thus, in two-player normal-form games, we can straightforwardly perform efficient opponent modeling and exploitation against static mixed strategies. This approach can also be straightforwardly extended to apply to more than two players whose strategies are played independently. In perfect-information sequential games, opponent modeling is also straightforward. Such games are typically represented as trees. We can simply apply the same procedure as above independently at each decision node of the tree. This also applies whether or not chance moves are allowed in the game tree.
@@
To perform opponent modeling we assume that we are in the repeated game setting where the game is repeated for some number of iterations. This number may be finite, infinite, or unknown (and potentially infinite). Defining repeated games in the normal-form setting is relatively straightforward. At each iteration $t$, we know our own mixed strategy $\sigma^{t}_{i},$ as well as the vector of pure strategies selected by all players $\mathbf{s^{t}}$ (though of course we do not know opponents’ mixed strategies). We remember all of these strategies for all subsequent iterations. Therefore, at iteration $t$, our strategy can be dependent on $\sigma^{1}_{i},\ldots,\sigma^{t-1}_{i}$ as well as $\mathbf{s^{1}},\ldots,\mathbf{s^{t-1}}$. In perfect-information sequential games, we know our own mixed strategy $\sigma^{t}_{i},$ as well as the actions taken by all players at nodes along the path of play (we do not know what an opposing player would have done at a node that was not reached). Again, at time $t$ our strategy can be dependent on these values for times from 1 to t1.t-1.
@@
Now that we have defined the repeated game setting for imperfect-information games, we can consider opponent modeling algorithms for this setting. It turns out that the general goal is still the same; given a prior distribution (possibly an uninformative Dirichlet distribution whose mean selects each available action with equal probability) and our observations made during gameplay, we would like our model for the opponent’s strategy to be the mean of the posterior distribution. This is justified by the following theorem and corollary [3]. Theorem 1 shows that the expected payoff of a mixed strategy $\sigma_{i}$ against a probability distribution with pdf $f_{-i}$ over mixed strategies equals the expected payoff of playing $\sigma_{i}$ against the mean $\overline{f_{-i}}$ of $f_{-i}$. This applies to both normal-form and extensive-form games (with perfect and imperfect information) as well as to any number of players ($\sigma_{-i}$ can be a joint strategy profile for all opposing agents). From this result Corollary 1 immediately follows if we let $f_{-i}$ be the posterior probability distribution for the opponent’s strategy $\sigma_{-i}$ given observations $x$, denoted as $p(\sigma_{-i}|x)$.
@@ 2 Consistent opponent modeling
If we can show that BBR is not consistent in a normal-form game, this directly implies that it is also not consistent in extensive-form games of perfect and imperfect information. Consider the game of Rock-Paper-Scissors and suppose that $\sigma^{*}_{-i}=(0.8,0.1,0.1)$ for (R,P,S). Suppose that we are using BBR with $k=3$ sampled strategies: (0.5,0.3,0.2), (0.3,0.5,0.2), (0.2,0.3,0.5). BBR chooses the opponent model to play each of the sampled strategies in proportion to their posterior value; therefore, the opponent model is always a convex combination of the sampled strategies. Since $\sigma^{*}_{-i}$ is clearly not close to any convex combination of the sampled strategies, clearly $M_{t}$ cannot approach $\sigma^{*}_{-i}$. ∎
As the proof of Proposition 1 shows, clearly BBR is constrained to only selecting opponent models that are convex combinations of the sampled strategies. So if the true strategy $\sigma^{*}_{-i}$ is outside the convex hull of the sampled strategies, BBR has no hope of the opponent model approaching the true strategy even if the game is repeated infinitely many times. Of course, in a normal-form game such as Rock-Paper-Scissors we could include all pure strategies of the game in our set of “samples,” which would ensure that every mixed strategy is within the convex hull of the samples. However, in extensive-form games there are exponentially many pure strategies in the size of the game tree and it is not feasible to include all of them. In fact, Proposition 2 shows that BBR may not be consistent even if $\sigma^{*}_{-i}$ is in the convex hull of the sampled strategies.
@@ 3 Algorithm
Proposition 4 (Consistency under standard assumptions).
Assume (i) the opponent’s true strategy $\sigma^{*}_{-i}$ lies in the model class and has positive prior probability, (ii) the data-generating process is correctly specified and identifies $\sigma^{*}_{-i}$ (i.e., distinct opponent strategies induce distinct distributions over the observations generated by the algorithm), and (iii) all opponent information sets/actions whose probabilities are being estimated are visited infinitely often (equivalently, the induced observation process provides persistent excitation). Then the algorithm is consistent: the posterior concentrates at $\sigma^{*}_{-i}$ and $\lim_{t\to\infty}M_{t}=\sigma^{*}_{-i}$.
@@
At each iteration $t$ the algorithm computes the modeled strategy $M_{t}$ that maximizes the posterior (MAP). Under assumptions (i)–(iii), standard Bayesian consistency results imply that the posterior distribution concentrates to a point mass at the true opponent strategy $\sigma^{*}_{-i}$ as the number of observations (iterations) goes to infinity. Therefore the posterior mode converges to $\sigma^{*}_{-i}$, i.e., $\lim_{t\rightarrow\infty}M_{t}=\sigma^{*}_{-i}$. ∎
Remark. If some opponent information sets are reached with zero probability under the (possibly exploitative) play induced by the algorithm, then the corresponding parameters are not identifiable from on-policy data and the conclusion can fail.
@@ 6 Conclusion
We introduced a new algorithm that computes the strategy for the opponent that maximizes the posterior distribution assuming independent Dirichlet prior distributions at each information set. The Dirichlet prior is very common for modeling vectors in a multidimensional simplex. With no further information available, we could choose a Dirichlet distribution with equal parameters for each action (whose mean selects each action with equal probability). If historical information is available, we could alternatively use a Dirichlet prior distribution whose mean strategy is calculated from the data. An additional option that has proven successful in certain applications is to use a Dirichlet prior whose mean strategy is a Nash equilibrium (or approximation of one) [2, 4]. Our algorithm runs efficiently by solving a convex minimization problem using projected gradient descent. We show that our algorithm achieves consistency under standard identifiability and visitation assumptions, unlike several prior algorithms that utilize sampling.